OpenStudy (anonymous):

Add 3 1/8 + 4 5/8 1 1/2 7 3/4 7 1/2 1 3/4

ganeshie8 (ganeshie8):

\[\huge 3\frac{1}{8} + 4\frac{5}{8}\]

ganeshie8 (ganeshie8):

add the integer part separately, then add the fraction parts.

OpenStudy (precal):

don't forget to reduce fractions at the end

ganeshie8 (ganeshie8):

\[\huge 3+4 + \frac{1}{8}+\frac{5}{8}\]

ganeshie8 (ganeshie8):

\[\huge 7\frac{6}{8}\] \[\huge = 7\frac{3}{4}\]